# Bayesian Controlled FDR Variable Selection via Knockoffs

### v1

@@ 5 Concluding remarks
In this article, we have proposed an approach that generalizes the model-X knockoff filter of Candès et al., (2018) to the Bayesian framework. Our formulation allows for variable selection while controlling the BFDR at an arbitrary level, making use of knockoffs as latent variables that act as negative controls. Additionally, we have imposed a sparse conditional independence structure on the covariates and used it in the prior on the knockoffs in a way that satisfies their pairwise conditional independence property. We have also limited the computational cost by adopting a modified version of the spike-and-slab prior that avoids the increase of the model dimension. Our model performs variable selection using an upper bound on the posterior probability of non-inclusion. We have shown how our model construction leads to valid model-X knockoffs and demonstrated that the proposed approach is sufficient for controlling the BFDR at an arbitrary level, in finite samples. We have also shown that the model selection is robust to the estimation of the precision matrix.
@@
In applications, we have shown that our proposal increases the stability of the selection, since it relies on the entire posterior distribution of the knockoff variables instead of a single sample. Results from our simulation study have shown how in simple settings (like Scenarios 1 and 2) our Bayesian knockoff filter approach has similar performance to state-of-art Bayesian variable selection techniques, such as spike-and-slab priors, while outperforming classical variable selection methods, including classical knockoffs filter. Moreover, under more complex dependence structure of the covariates (as in Scenario 3), our proposal is able to control the FDR under any desired threshold, unlike spike-and-slab variable selection methods, that suffer from poor estimates of the PPIs. It also performs as well as classical knockoff filter with covariance structure known. This is relevant since in practice the precision matrix is never known. As an illustrative example, we have applied the model to a benchmark data set on prostate cancer collected from the study of Stamey et al., (1989), and found more relevant variables than the classical knockoff filter.
@@
Future work could extend this research in several directions. First, while we have assumed a standard Gaussian linear regression setting, similar Bayesian knockoff filters could be designed for linear settings that incorporate discrete-type responses. For example, in the simulations we have also considered an AFT model for survival outcomes. A similar extension, that uses data augmentation techniques, can be used for multinomial probit models (Albert and Chib,, 1993; Sha et al.,, 2004). Furthermore, logit and negative binomial responses could be accommodated via the use of Polya-Gamma data augmentation schemes (Polson et al.,, 2013; Zhou et al.,, 2012). Another interesting direction would be to extend the proposed knockoffs framework to the case of non-Gaussian covariates. The generation of knockoffs for non Gaussian random variables is still an open research problem. This case presents the main challenge of not having a known distribution that respects pairwise exchangeability (Bates et al.,, 2021). A promising direction is represented by the approach proposed by Dreassi et al., (2024), which makes use of conditional independence to build the knockoffs variables. Finally, the variable selection prior distribution we have used on the coefficients presents some computational difficulties due to the need to use a Metropolis-within-Gibbs algorithm. As an alternative, a shrinkage prior such as the horseshoe prior (Carvalho et al.,, 2010), could be used to reduce computational costs. However, this reduced computational cost comes at the expense of greater difficulty in utilizing the information from the estimated graph on the covariates.
Appendix A Complete proof of Proposition 1
Theorem A.1.
Given the model in Equation (9), let $p(\beta,\widetilde{\beta})$ be a prior distribution such that
$p(\beta,\widetilde{\beta})_{\text{Swap}(j)}=p(\beta,\widetilde{\beta})\qquad j=1,\dots,p$
and let $W=(W_{1},\dots,W_{p})^{T}$ be an anti-symmetric function of $\beta$ and $\widetilde{\beta}$. Let $r=(r_{1},\dots,r_{p})\in\{0,1\}^{p}$ be a random vector such that $r_{j}=0\iff X_{j}\mbox{$\perp\!\!\!\perp\,$}Y\mid X_{-j}$. Then the posterior distribution of $W$ obeys the sign-flip property, i.e.
$\mathbb{P}[W_{j}>0\mid\bm{D},r_{j}=0]=\mathbb{P}[W_{j}<0\mid\bm{D},r_{j}=0]$
Proof.
In the context of this proof, the model in Equation (9) is equivalent to the simple linear model reported in Equation (7). The marginalization of latent knockoffs is, indeed, not required since for this proof both the conditional independence and the pairwise exchangeability properties are only required to be true a priori. Moreover, applying the knockoff decomposition in Equation (4) and working with their random part $U$ is equivalent to work with the complete knockoffs $\widetilde{X}$. Thus, we will rely on the complete knockoffs for simplicity. Therefore, this proof, which we report for completeness, follows from results in Gu and Yin, (2021)
For all $j$ such that $r_{j}=0$, it holds that $X_{j}\mbox{$\perp\!\!\!\perp\,$}Y|X_{-j}$. Thus, we can write $X_{\mathcal{S}}\mbox{$\perp\!\!\!\perp\,$}Y|X_{-\mathcal{S}}$ where the set $\mathcal{S}=\{j\in\{1,\dots,p\}:r_{j}=0\}$. Moreover, by prior conditional independence, we have
$X_{\mathcal{S}}\mbox{$\perp\!\!\!\perp\,$}Y|X_{-\mathcal{S}},\widetilde{X}.$
Since $p(X,\widetilde{X})_{\text{swap}(j)}=p(X,\widetilde{X})$, it is also true that $\widetilde{X}_{\mathcal{S}}\mbox{$\perp\!\!\!\perp\,$}Y|\widetilde{X}_{-\mathcal{S}},X$. This implies
$p(Y|X,\widetilde{X})=p(Y|X_{\text{swap}(\mathcal{S})},\widetilde{X}_{\text{swap}(\mathcal{S})}).$
Equation (7) implies that the response variable and the augmented covariate matrix $[X,\widetilde{X}]$ are related only through regression coefficient. So
$p(Y|X_{\text{swap}(\mathcal{S})},\widetilde{X}_{\text{swap}(\mathcal{S})};\beta,\widetilde{\beta},\sigma)=p(Y|X,\widetilde{X};\beta_{\text{swap}(\mathcal{S})},\widetilde{\beta}_{\text{swap}(\mathcal{S})},\sigma).$
Since the prior probability distribution of the coefficient is invariant to swaps, it holds that
$p(\beta,\widetilde{\beta}|\widetilde{x}_{1},\dots,\widetilde{x}_{n},\sigma,\bm{D})=p(\beta_{\text{swap}(\mathcal{S})},\widetilde{\beta}_{\text{swap}(\mathcal{S})}|\widetilde{x}_{1},\dots,\widetilde{x}_{n},\sigma,\bm{D}).$
We can, then, write the marginal posterior distribution of the coefficients as
[…]
$=p(\beta_{\text{swap}(\mathcal{S})},\widetilde{\beta}_{\text{swap}(\mathcal{S})}|\bm{D}).$
Define the vector $W_{\mathcal{S}}=(W_{1,\mathcal{S}},\dots,W_{p,\mathcal{S}})$ as
$W_{j,\mathcal{S}}=\begin{cases}W_{j}&r_{j}=1\\ -W_{j}&r_{j}=0.\end{cases}$
Finally, this implies the following
[…]
$=p(W|\bm{D})$
where $\Lambda_{W}$ is the subspace of values of the vector $(\beta,\widetilde{\beta})$ that share the same value of $W$. Note that the second equality follows by the fact that $W$ is an antisymmetric function. ∎
Appendix B Detailed MCMC procedure
Here we provide further details on the MCMC procedure. Overall, our proposed algorithm can be summarized by the following steps:
  • (a)
    Update the set of variable selection parameters ($\delta$, $\widetilde{\delta}$, $\gamma$) and the regression coefficients ($\beta$, $\widetilde{\beta}$). Since the regression coefficients $\beta$ and $\widetilde{\beta}$ are necessary to the construction of the statistics $W$ used for variable selection, we cannot integrate them out. Therefore, we propose a joint update of the parameters based upon randomly sampling one index $i\in\{1,\dots,p\}$ and changing the value of the variable inclusion indicator $\gamma_{i}$. Once the proposed value $\gamma^{*}$ is obtained, we assign values for the parameters $\delta$ and $\widetilde{\delta}$. Specifically, if $\gamma^{*}_{j}=0$ both $\delta^{*}_{j}$ and $\widetilde{\delta}^{*}_{j}$ are set to zero. Otherwise, we randomly set $\delta_{j}^{*}$ or $\widetilde{\delta}^{*}_{j}$ to $1$. Finally, the proposed value for the regression coefficient is drawn from a Gaussian distribution with mean 00 and variance $0.5$ times the corresponding proposed indicator $\delta^{*}_{j}$ and $\widetilde{\delta}^{*}_{j}$ for $\beta^{*}$ and $\widetilde{\beta}^{*}_{j}$, respectively. Note that the moves induced by the proposal density $q(\gamma,\delta,\widetilde{\delta},\beta,\widetilde{\beta})$ are symmetric therefore the proposal ratio is equal to 1 and and the Metropolis-Hastings ratio depends only on the likelihood ratio and the ratio of the priors. We can write the likelihood ratio $r_{Y}$ as
    $r_{Y}$ $=\frac{p(Y|X,U,\beta^{*},\widetilde{\beta}^{*},\bm{\Omega},\sigma^{2})}{p(Y|X,U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})}$
    $=\exp\bigg{\{}-\frac{1}{2\sigma^{2}}\bigg{[}(y-\bm{X}[\beta^{*}+(I-\bm{\Omega}\text{diag}(s))\widetilde{\beta}^{*}]-U\widetilde{\beta}^{*})^{T}(y-\bm{X}[\beta^{*}+(I-\bm{\Omega}\text{diag}(s))\widetilde{\beta}^{*}]$
    $-U\widetilde{\beta}^{*})-(y-\bm{X}[\beta+(I-\bm{\Omega}\text{diag}(s))\widetilde{\beta}]-U\widetilde{\beta})^{T}(y-\bm{X}[\beta+(I-\bm{\Omega}\text{diag}(s))\widetilde{\beta}]-U\widetilde{\beta})\bigg{]}\bigg{\}}.$
    Thus, the acceptance ratio of the Metropolis-Hastings algorithm of step (a) is
    $r$ $=r_{Y}\times\frac{p(\beta^{*},\widetilde{\beta}^{*}\mid\gamma^{*},\sigma^{2},h_{\beta})p(\delta^{*},\widetilde{\delta}^{*}\mid\gamma^{*})p(\gamma^{*}\mid\bm{G},a,b)}{p(\beta,\widetilde{\beta}\mid\gamma,\sigma^{2},h_{\beta})p(\delta,\widetilde{\delta}\mid\gamma)p(\gamma\mid\bm{G},a,b)}$
    $=r_{Y}\times\exp\{a\mathbbm{1}(\gamma^{*}-\gamma)+b(\gamma^{*}-\gamma)^{T}\bm{G}(\gamma^{*}+\gamma)\}\times r_{\beta}\times r_{\delta},$
    where
    $r_{\delta}=\begin{cases}0.5&\text{if}\quad\gamma^{*}_{j}=1\\ 2&\text{if}\quad\gamma_{j}^{*}=0\end{cases}$
    $r_{\beta}=\begin{cases}\frac{1}{2\pi h_{\beta}\sigma^{2}}\exp\left\{-\frac{1}{2\sqrt{h_{\beta}}\sigma}\beta_{j}^{*^{2}}\right\}&\text{if}\quad\gamma^{*}_{j}=1\text{ and }\delta^{*}_{j}=1\\ \frac{1}{2\pi h_{\beta}\sigma^{2}}\exp\left\{-\frac{1}{2\sqrt{h_{\beta}}\sigma}\widetilde{\beta}_{j}^{*^{2}}\right\}&\text{if}\quad\gamma^{*}_{j}=1\text{ and }\widetilde{\delta}^{*}_{j}=1\\ 2\pi h_{\beta}\sigma^{2}\exp\left\{\frac{1}{2\sqrt{h_{\beta}}\sigma}\beta_{j}^{{}^{2}}\right\}&\text{if}\quad\gamma^{*}_{j}=0\text{ and }\delta_{j}=1\\ 2\pi h_{\beta}\sigma^{2}\exp\left\{\frac{1}{2\sqrt{h_{\beta}}\sigma}\widetilde{\beta}_{j}^{{}^{2}}\right\}&\text{if}\quad\gamma^{*}_{j}=0\text{ and }\widetilde{\delta}_{j}=1\\ \end{cases}.$
    To improve the mixing of the algorithm, step (a) contains a within-model update where for each index $i$ such that $\delta_{i}=1$ and for each $j$ such that $\widetilde{\delta}_{j}=1$, we propose a new value for the corresponding regression coefficient $\beta_{i}$ and $\widetilde{\beta}_{j}$. In this step, the proposal distribution $q(\beta)$ is a Gaussian distribution with mean the current value of $\beta$ and variance $0.5$. To simplify notation, we focus on the within-model update for $\beta$ since it is exactly the same for $\widetilde{\beta}$. The acceptance ratio for this step $r_{W}$ is
    $r_{W}=r_{Y}\times\exp\left\{-\frac{1}{2\sigma\sqrt{h_{\beta}}}\left(\beta_{i}^{*^{2}}-\beta_{i}^{2}\right)\right\}.$
  • (b)
    Update the graph parameters ($\bm{G}$, $\bm{\Omega}$). This is done by Gibbs sampling using the full conditional distributions via the efficient block-sampling algorithm proposed by Wang, (2015). First, each column of the precision matrix $\Omega$ is sampled from its full conditional distribution. To do so, consider the following change of variable $(\omega_{12},\omega_{22})\mapsto(k=\omega_{12},\,l=\omega_{22}-\omega_{12}^{T}\bm{\Omega}_{11}^{-1}\omega_{12})$ where the precision matrix has been decomposed as
    $\bm{\Omega}=\begin{pmatrix}\bm{\Omega}_{11}&\omega_{12}\\ \omega_{12}^{T}&\omega_{22}\end{pmatrix}.$
    Then, Wang, (2015) showed how the full conditional of $k$ is a Gaussian distribution with mean $-\bm{C}s_{12}$ and covariance matrix $\bm{C}$ where $\bm{C}=\left\{(s_{22}+\theta)\bm{\Omega}_{11}^{-1}+\text{diag}(1/v_{12})\right\}^{-1}$. The full conditional of $l$ is, instead, a Gamma distribution with parameters $n/2+1$ and $(s_{22}+\theta)/2$. In the full conditional distributions, $\bm{S}$ is the sample covariance matrix, and $\bm{V}$ is a matrix with 00 on the diagonal and $v_{g_{jj^{\prime}}}$ on the off-diagonal entries. Both of these matrices have been decomposed as $\bm{\Omega}$. Once the precision matrix $\bm{\Omega}$ has been updated, the full conditional distribution of the adjacency matrix is
    $\mathbb{P}(g_{i,j}=1\mid\bm{\Omega})=\frac{N(\omega_{ij}\mid 0,v^{2}_{1})\xi}{N(\omega_{ij}\mid 0,v^{2}_{1})\xi+N(\omega_{ij}\mid 0,v^{2}_{0})(1-\xi)}.$
  • (c)
    Update the knockoff latent variables ($U$). This follows directly from a basic normal-normal model. We, therefore, omit the completing the square procedure. The resulting full conditional distributions of each row of the latent variable $U$ are reported in Section 2.2.3.
In the case of the AFT model used in the simulations, it is possible to use the exact same sampling scheme of a linear Gaussian model with an added step to update the value of the latent $y_{i}$. This update is straightforward since we do not integrate out any parameter of the model. Specifically, for subjects with $\kappa_{i}=0$, the full conditional of $y_{i}$ is the following Gaussian distribution
$y_{i}\mid X,U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2}\sim N(X_{i}[\beta+(I-\bm{\Omega}\text{diag}(s))\widetilde{\beta}]+U_{i}\widetilde{\beta},\sigma^{2})$
truncated in the interval $(c_{i},\infty)$.

### v2

@@ 5 Concluding remarks
In this article, we have proposed a fully Bayesian generalization of the model-X knockoff filter of Candès et al., (2018) that allows for variable selection while controlling the BFDR at an arbitrary level. Our framework treats knockoff variables as latent variables that act as negative controls. Additionally, we have imposed a sparse conditional independence structure on the covariates and used it in the prior on the knockoffs in a way that satisfies their pairwise conditional independence property. We have also limited the computational cost by adopting a modified version of the spike-and-slab prior that avoids the increase of the model dimension. Our model performs variable selection using an upper bound on the posterior probability of non-inclusion. We have shown how our model construction leads to valid model-X knockoffs and demonstrated that the proposed approach is sufficient for controlling the BFDR at an arbitrary level, in finite samples, if the distribution of the covariates is fully known, and asymptotically if estimated as in our model.
@@
In applications, we have shown that our proposal increases the stability of the selection, since it relies on the entire posterior distribution of the knockoff variables instead of a single sample. Results from our simulation study have shown how in simple settings (like Scenarios 1 and 2) our Bayesian knockoff filter approach has similar performance to state-of-art Bayesian variable selection techniques, such as spike-and-slab priors, while outperforming classical variable selection methods, including classical knockoffs filter. Moreover, under more complex dependence structure of the covariates (as in Scenario 3), our proposal is able to control the FDR under any desired threshold, unlike spike-and-slab variable selection methods, that suffer from poor estimates of the PPIs. It also performs as well as classical knockoff filter with covariance structure known. This is relevant since in practice the precision matrix is never known.
@@
Future work could extend this research in several directions. First, while we have assumed a standard Gaussian linear regression setting, similar Bayesian knockoff filters could be designed for linear settings that incorporate discrete-type responses. For example, in the simulations we have also considered an AFT model for survival outcomes. A similar extension, that uses data augmentation techniques, can be used for multinomial probit models (Albert and Chib,, 1993; Sha et al.,, 2004). Furthermore, logit and negative binomial responses could be accommodated via the use of Polya-Gamma data augmentation schemes (Polson et al.,, 2013; Zhou et al.,, 2012). Another interesting direction would be to extend the proposed knockoffs framework to the case of non-Gaussian covariates. The generation of knockoffs for non Gaussian random variables is still an open research problem. This case presents the main challenge of not having a known distribution that respects pairwise exchangeability (Bates et al.,, 2021). A promising direction is represented by the approach proposed by Dreassi et al., (2024), which makes use of conditional independence to build the knockoffs variables. Finally, the variable selection prior distribution we have used on the coefficients presents some computational challenges due to the need to use a Metropolis-within-Gibbs algorithm. Since the largest computational burden comes from the generation of the latent knockoffs at each MCMC iteration, a possible solution would be to use variational inference (Blei et al.,, 2017). However, the restrictive assumptions to obtain valid knockoffs require careful design of the approximate variational distributions. In particular, the commonly used mean field approximation cannot be used since it breaks the pairwise exchangeability assumption of the knockoffs and the covariates.